OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 2y = 6 x + 3y = −1 a. -2 b. 2 c. 5 d. -5

OpenStudy (allank):

I like using elimination method. Multiply the second equation by 2: 2(x+3y)=2(-1) 2x+6y=-2 Then subtract the second equation from the first. (2x+2y=6)-(2x+6y=-2) That eliminates the 2x to get -4y=8 y=-2 To get x, plug in y=-2 in the first or second equations. I choose the second. x + 3y = −1 x=-1-3y x=-1-3(-2) x=-1+6 x=5